micheal received $275 more from relatives than friends. he receives $4275 dollars in all, how much more money did he recieve from relatives and how much dis he receive from friends ?
Answer by 4198181(13)

Let r represent the amount of money from relatives
Let f represent the amount of money from friends
f+r=4275
f+275=r
Use substitution, f+(f+275)=4275
2f+275=4275
2f=4000
f=2000
Substitute 2000 in for f
2000+275=r
2275=r
ANSWER:
Micheal received $2275 from relatives and $2000 from friends